# Random Lévy Looptrees and Lévy Maps

### v2

@@ 4.1.   Hölder continuity of the distances
Recall from (1.2) the upper and lower exponents $\boldsymbol{\eta}$ and $\boldsymbol{\gamma}$ of $\psi$ at infinity. The main goal of this subsection is to prove that the canonical projections $[0,1]\to\mathscr{L}(X^{\mathrm{ex}})$ and $[0,1]\to M$ are almost surely Hölder continuous with exponent $a$ and $a/2$ respectively for any $a\in(0,1/\boldsymbol{\eta})$. We shall need the following tail bound.
@@
There exists a function $h\colon(0,\infty)\to(0,\infty)$ and a constant $C>0$ such that for every $t,x>0$,
$\mathop{}\!\mathbb{P}\Bigl{(}\sup_{s\in[0,t]}|X_{s}|>x\Bigr{)}\leq C\cdot t\cdot h(x).$
Furthermore, it has the following asymptotic behaviour near 00:
$\inf\Bigl{\{}c\geq 0\colon\limsup_{x\downarrow 0}x^{c}h(x)=0\Bigr{\}}=\boldsymbol{\eta}\qquad\text{and}\qquad\inf\Bigl{\{}c\geq 0\colon\liminf_{x\downarrow 0}x^{c}h(x)=0\Bigr{\}}=\boldsymbol{\gamma}.$ (4.1)
@@
Let us first assume that $X$ has no drift and no Gaussian component. We then deduce the general case in the end. The bound on the tail probability is given by Pruitt [Pru81], even in the case where $X$ may also have negative jumps, see Equation 3.2 there. By combining [Pru81, Equation 3.4] with the work [BG61], one can derive the upper exponent $\boldsymbol{\eta}$ for $h$, but only lower bound the lower exponent of $h$ by $\boldsymbol{\gamma}$. Let us give an alternative proof of and prove that $\boldsymbol{\gamma}$ is precisely the lower exponent of $h$. The function $h$ is given in [Pru81] (up to the choice of the convention for the Lévy–Khintchine formula) by:
$h(x)=\overline{\pi}(x)+x^{-2}\int_{(0,x]}r^{2}\pi(\mathop{}\!\mathrm{d}r)+x^{-1}\int_{(x,\infty)}r\pi(\mathop{}\!\mathrm{d}r)$
where we have set $\overline{\pi}(x)=\pi((x,\infty))$. We claim that
$h(x)\asymp\psi(1/x)\qquad\text{in }(0,1),$ (4.2)
by which we mean that there exist two constants $c_{1},c_{2}>0$ such that $c_{1}\psi(1/x)\leq h(x)\leq c_{2}\psi(1/x)$ for every $x\in(0,1)$. The exponents of $h$ will then follow from those of $\psi$ in (1.2).
To establish (4.2) we rely on [Duq12, Section 2.3]. Indeed, recall that $\phi=\psi^{\prime}$ is the Laplace exponent of a subordinator which has Lévy measure $\nu(\mathop{}\!\mathrm{d}r)=r\pi(\mathop{}\!\mathrm{d}r)$. Then $h(x)-\overline{\pi}(x)$ equals precisely $x^{-2}J_{\psi^{\prime}}(x)$ in the notation of [Duq12, Equation 47]. Recall the notation $\widetilde{\psi}(u)=\psi(u)/u$. By combining Equations 47 and 50 in [Duq12], we infer that
$h(x)-\overline{\pi}(x)=x^{-2}J_{\psi^{\prime}}(x)\asymp x^{-1}\psi^{\prime}(1/x)\asymp x^{-1}\widetilde{\psi}(1/x)=\psi(1/x),$
in $(0,1)$. This already gives the lower bound in (4.2) by simply lower bounding $\overline{\pi}$ by 00. As for the upper bound, simply observe that for every $x>0$, we have $\overline{\pi}(x)\leq x^{-1}\int_{(x,\infty)}r\pi(\mathop{}\!\mathrm{d}r)$ so $\overline{\pi}$ is also upper bounded by some constant times $\psi(1/\cdot)$ by the previous argument. This concludes the proof of of (4.2) when $\beta=0$.
In the case where $X$ has a drift $d\neq 0$ and a Gaussian parameter $\beta>0$, the Lévy–Itō decomposition reads $X=D+W+Y$ in distribution, where $D_{t}=dt$, where $W$ is $\sqrt{\beta}$ times a standard Brownian motion, and where $Y$ is independent of $W$ and has no drift nor Gaussian component. Recall that in this case, we have $\psi(\lambda)\sim\beta\lambda^{2}$ as $\lambda\to\infty$. By the first part of the proof, $Y$ satisfies the claim with some $1\leq\boldsymbol{\gamma}\leq\boldsymbol{\eta}\leq 2$. So does $W$, with $h(x)=x^{-2}$, by e.g. an application of the reflection principle and a well-known tail bound for Gaussian random variables:
$\mathop{}\!\mathbb{P}\Bigl{(}\sup_{s\in[0,t]}|W_{s}|>x\Bigr{)}\leq 4\mathop{}\!\mathbb{P}(W_{t}>x)\leq 4\exp\Bigl{(}-\frac{x^{2}}{2\beta t}\Bigr{)}\leq\frac{4}{\mathrm{e}}\frac{2\beta t}{x^{2}}.$
The claim, with $h(x)=x^{-2}$, then easily follows from a union bound. ∎
Remark 4.2.
Let us mention that the lower exponent of $\overline{\pi}$ near 00 may be smaller than $\boldsymbol{\gamma}$, even in the case of a subordinator, see e.g. Theorems 6.1 and 7.1 in [BG61] for some more information.
Recall from [Mar22b, Proposition 3.2] that the distance function $d_{\mathscr{L}(X^{\mathrm{ex}})}$ from (2.6) is almost surely a continuous pseudo-distance, so the canonical projection $\mathbf{p}_{\mathscr{L}(X^{\mathrm{ex}})}\colon[0,1]\to\mathscr{L}(X^{\mathrm{ex}})$ is almost surely continuous. The next proposition shows that it is Hölder continuous; we will easily derive from this the upper bound for the Hausdorff and upper Minkowski dimensions of the looptree.
Proposition 4.3.
For any $a\in(0,1/\boldsymbol{\eta})$, the projection $\mathbf{p}_{\mathscr{L}(X^{\mathrm{ex}})}$ is almost surely Hölder continuous with exponent $a$ in the sense that almost surely, there exists $C_{a}\in(0,\infty)$ such that:
We shall see that the exponent $1/\boldsymbol{\eta}$ is optimal in the sense that $\mathbf{p}_{\mathscr{L}(X^{\mathrm{ex}})}$ is almost surely not Hölder continuous with exponent larger than $1/\boldsymbol{\eta}$, see Remark 4.7. As the proof of will show, the claim also holds for the unconditioned Lévy process $X$ and its bridge version.
@@
For $n\leq n_{0}$, there are only finitely many quantities involved, so we infer that almost surely, there exists a now random $C_{a}\in(0,\infty)$ such that (4.3) holds for every pair $(s,t)$ of the form $(j2^{-n},(j+1)2^{-n})$ with $n\geq 1$ and $j\in\{0,\dots,2^{n}-1\}$. It then extends to all dyadic numbers by a routing argument, after multiplying $C_{a}$ by $2/(1-2^{-a})$, and then to the whole interval $[0,1]$ by density, using that $d_{\mathscr{L}(X^{\mathrm{ex}})}$ is continuous almost surely. It remains to prove (4.4). Let us first replace $X^{\mathrm{ex}}$ by the unconditioned process $X$ and then argue by absolute continuity. According to (2.7), almost surely, for $s<t$, we have:
$d_{\mathscr{L}(X^{\mathrm{ex}})}(s,t)\leq X_{s}+X_{t-}-2\inf_{r\in[s,t]}X_{r}\leq 4\sup_{r\in[s,t]}|X_{r}-X_{s}|.$
@@
Indeed, informally the term in the middle is the length of the path if one goes from $s$ to $t$ in the looptree by always following a cycle on its right, instead of optimising between left and right. Note that the Markov property implies that the right-hand side has the same law as $4\sup_{r\in[0,t-s]}|X_{r}|$. This bound is too crude to be used directly because of the jumps of $X$, but we shall use it between two consecutive large jumps.
@@
Precisely, fix $\delta>0$ and let $\theta=(1+\delta)(\boldsymbol{\eta}+\delta)$; we shall take $a=1/\theta$. Fix $n\geq 1$ and $i\in\{0,\dots,2^{n}-1\}$. Let $(t_{n,i}(1),\dots,t_{n,i}(N_{n,i}))$ denote the increasing enumeration of the times $\{t\in[i2^{-n},(i+1)2^{-n}]\colon\Delta X_{t}>2^{-n/\theta}\}$. Set also $t_{n,i}(0)=i2^{-n}$ and $t_{n,i}(N_{n,i}+1)=(i+1)2^{-n}$. Then combining the last display with the triangle inequality, we have:
$d_{\mathscr{L}(X^{\mathrm{ex}})}(i2^{-n},(i+1)2^{-n})\leq 4\sum_{j=1}^{N_{n,i}}\sup_{r\in[t_{n,i}(j),t_{n,i}(j+1)]}|X_{r}-X_{t_{n,i}(j)}|.$ (4.5)
Let us first bound $N_{n,i}$. The latter has the Poisson distribution with rate $2^{-n}\overline{\pi}(2^{-n/\theta})$ and the Chernoff bound shows that the tail probability of a Poisson random variable $Y$ with rate $\lambda$ satisfies $\mathop{}\!\mathbb{P}(Y\geq k)\leq\operatorname{e}^{-\lambda}(\operatorname{e}\lambda/k)^{k}$ for every $k>\lambda$. Hence for any integer $N\geq 1$, there exists a constant $C_{N}$, which depends on $N$, such that:
$\mathop{}\!\mathbb{P}(N_{n,i}\geq N)\leq C_{N}(2^{-n}\overline{\pi}(2^{-n/\theta}))^{N}.$
Now recall from the proof of Lemma 4.1 that $\overline{\pi}$ is upper bounded by some constant times $\psi(1/\cdot)$, so there exists a constant $C_{\delta}>0$ such that $\overline{\pi}(x)\leq C_{\delta}x^{-(\boldsymbol{\eta}+\delta)}=C_{\delta}x^{-\theta/(1+\delta)}$ for every $x\in(0,1)$. We infer that
$\mathop{}\!\mathbb{P}(N_{n,i}\geq N)\leq C_{N}C_{\delta}^{N}2^{-nN\delta/(1+\delta)}.$
Let us next bound the variation of $X$ between $t_{n,i}(j)$ and $t_{n,i}(j+1)$. By definition, between these times, the process $X$ makes no jump larger than $2^{-n/\theta}$, therefore, at the first passage time above this level, the value is at most $2\times 2^{-n/\theta}$. Using the strong Markov property, one then easily gets by induction that for every integer $M\geq 1$, it holds:
$\mathop{}\!\mathbb{P}\Bigl{(}\sup_{r\in[0,t_{n,i}(j+1)-t_{n,i}(j)]}|X_{r}|>2M2^{-n/\theta}\bigm{|}t_{n,i}(j),t_{n,i}(j+1)\Bigr{)}\leq\mathop{}\!\mathbb{P}\Bigl{(}\sup_{r\in[0,t_{n,i}(j+1)-t_{n,i}(j)]}|X_{r}|>2^{-n/\theta}\bigm{|}t_{n,i}(j),t_{n,i}(j+1)\Bigr{)}^{M}.$
We next upper bound the probability on the right thanks to Lemma 4.1. There exists a constant $C_{\delta}>0$ such that, with the constant $C$ and the function $h$ from this lemma, we have $Ch(x)\leq C_{\delta}x^{-(\boldsymbol{\eta}+\delta)}=C_{\delta}x^{-\theta/(1+\delta)}$ for every $x\in(0,1)$. Hence
$\mathop{}\!\mathbb{P}\Bigl{(}\sup_{r\in[0,t_{n,i}(j+1)-t_{n,i}(j)]}|X_{r}|>2^{-n/\theta}\bigm{|}t_{n,i}(j),t_{n,i}(j+1)\Bigr{)}\leq C2^{-n}h(2^{-n/\theta})\leq C_{\delta}2^{-n\delta/(1+\delta)}.$
Combined with union bounds, we obtain:
$\mathop{}\!\mathbb{P}\Bigl{(}N_{n,i}<N\quad\text{and}\quad 4\sum_{j=0}^{N_{n,i}}\sup_{r\in[t_{n,i}(j),t_{n,i}(j+1)]}|X_{r}-X_{t_{n,i}(j)}|>8N^{2}2^{-n/\theta}\Bigr{)}$
$\qquad\leq\mathop{}\!\mathbb{P}\Bigl{(}N_{n,i}<N\quad\text{and}\quad\exists j\leq N_{n,i}\colon\sup_{r\in[t_{n,i}(j),t_{n,i}(j+1)]}|X_{r}-X_{t_{n,i}(j)}|>2N2^{-n/\theta}\Bigr{)}$
$\qquad\leq NC_{\delta}^{M}2^{-nN\delta/(1+\delta)}.$
We have thus shown that for every $N\geq 1$, there exists a constant $K$ which depends on $N$ and $\delta$, but neither on $i$ nor $n$, such that
$\mathop{}\!\mathbb{P}(d_{\mathscr{L}(X^{\mathrm{ex}})}(i2^{-n},(i+1)2^{-n})>8N^{2}2^{-n/\theta})\leq K2^{-nN\delta/(1+\delta)}.$
A further union bound implies then:
$\mathop{}\!\mathbb{P}(\exists i\leq 2^{n}-1\colon d_{\mathscr{L}(X^{\mathrm{ex}})}(i2^{-n},(i+1)2^{-n})>8N^{2}2^{-n/\theta})\leq K2^{n-nN\delta/(1+\delta)}.$
If we take $N$ such that $N\geq 2+\frac{1}{\delta}$, then the sum over all $n\geq 0$ of the right-hand side is finite; the Borel–Cantelli lemma then yields (4.4) for the unconditioned Lévy process.
@@
Actually, more than (4.4) we have shown that its upper bound (4.5), which is expressed in a simple way in terms of the Lévy process, is itself bounded as in (4.4). This immediately extends to the first half the bridge $(X^{\mathrm{br}}_{t};0\leq t\leq 1/2)$, taking only the supremum over $i\in\{0,\dots,2^{n-1}\}$, by absolutely continuity with respect to the unconditioned process $(X_{t};0\leq t\leq 1/2)$. In addition, the time- and space-reversal of the second half of the bridge has the same law as the first half, so the upper bound also holds for the second half and thus for the entire bridge. Finally we transfer it to the excursion by applying the Vervaat transform. Notice that the latter changes the starting point and thus shifts the intervals, but this only mildly affects the constant $C_{a}$. ∎
@@
Note that if one wanted to replace the exponent $1/\boldsymbol{\eta}$ by the smaller one $1/\boldsymbol{\gamma}$, then instead of the uniform upper bound $h(x)\leq C_{\delta}x^{-(\boldsymbol{\eta}+\delta)}$ for every $x\in(0,1)$, one would only have the existence of a sequence $x_{n}\downarrow 0$ along which $\overline{\pi}(x_{n})\leq h(x_{n})\leq C_{\delta}x_{n}^{-(\boldsymbol{\gamma}+\delta)}$. Taking a subsequence if necessary, we may assume that $\sum_{n}x_{n}<\infty$ and then the preceding argument shows that instead of (4.4) for any $a\in(0,1/\boldsymbol{\gamma})$, almost surely there exist $K_{a}>0$ and a sequence $\varepsilon_{n}\downarrow 0$ such that, for every $n$ large enough,
@@
We then turn to the map $M$ and the label process $Z^{\mathrm{ex}}$. Informally, we simply loose a factor $1/2$ due to the regularity of the Brownian labels.
@@
As for looptree distances, the claim also holds for the label process associated with the unconditioned Lévy process or its bridge version.

### v3

@@ 5.1.   Hölder continuity of the distances
The main goal of this subsection is to prove that the canonical projections $[0,1]\to\mathscr{L}(X^{\mathrm{ex}})$ and $[0,1]\to M$ are almost surely Hölder continuous. We will easily derive from this the upper bound for the Hausdorff and upper Minkowski dimensions. Let us start with the looptree. Recall from (1.2) the upper and lower exponents $\boldsymbol{\eta}$ and $\boldsymbol{\gamma}$ of $\psi$ at infinity.
Proposition 5.1.
For any $a\in(0,1/\boldsymbol{\eta})$, there exists $C_{a}\in(0,\infty)$ such that:
We shall see that the exponent $1/\boldsymbol{\eta}$ is optimal in the sense that almost surely, no such $C_{a}<\infty$ exist for $a>1/\boldsymbol{\eta}$, see Remark 5.7. As the proof will show, the claim also holds for the unconditioned Lévy process $X$ and its bridge version.
We shall need the following tail bound that we state as a separate lemma. The ‘universal constant’ does not depend on the Lévy process.
@@
There exists a universal constant $C$ such that the following holds. For every Lévy process with no negative jump, for any $t,x>0$, we have
$\mathop{}\!\mathbb{P}\Bigl{(}\sup_{s\in[0,t]}|X_{s}|>x\Bigr{)}+t\,\overline{\pi}(x)\leq C\,t\,(2d^{+}x^{-1}+\psi(x^{-1})),$
where $d^{+}=\max(d,0)$ is the positive part of the drift and $\overline{\pi}(x)=\pi((x,\infty))$ is the tail of the Lévy measure.
@@
Let us focus on the probability on the left. Such a maximal inequality has been obtained by Pruitt [Pru81] for Lévy processes with no Gaussian component (but allowing for both positive and negative jumps). Precisely, Equation 3.2 there reads, when $\beta=0$:
$\mathop{}\!\mathbb{P}\Bigl{(}\sup_{s\in[0,t]}|X_{s}|>x\Bigr{)}\leq C\,t\,h(x),$
where, taking note of the difference in the convention for the Lévy–Khintchine formula,
$h(x)$ $=\overline{\pi}(x)+x^{-2}\int_{(0,x]}r^{2}\pi(\mathop{}\!\mathrm{d}r)+x^{-1}\Bigl{|}d-\int_{(x,\infty)}r\pi(\mathop{}\!\mathrm{d}r)\Bigr{|}$
$\leq\overline{\pi}(x)+x^{-2}\int_{(0,x]}r^{2}\pi(\mathop{}\!\mathrm{d}r)+x^{-1}\int_{(x,\infty)}r\pi(\mathop{}\!\mathrm{d}r)+|d|x^{-1}$
$\leq 2\Bigl{(}x^{-2}\int_{(0,x]}r^{2}\pi(\mathop{}\!\mathrm{d}r)+x^{-1}\int_{(x,\infty)}r\pi(\mathop{}\!\mathrm{d}r)+|d|x^{-1}\Bigr{)},$
after noticing $\overline{\pi}(x)\leq x^{-1}\int_{(x,\infty)}r\pi(\mathop{}\!\mathrm{d}r)$.
We may add the Gaussian component by recalling the identity $X=\beta W+Y$ in distribution, where $W$ is a standard Brownian motion, and where $Y$ is independent of $W$ and has no Gaussian component. By the reflection principle, a well-known tail bound for Gaussian random variables, and finally the easy bound $\operatorname{e}^{u}\geq\mathrm{e}\times u$ for every $u\geq 0$, we have:
$\mathop{}\!\mathbb{P}\Bigl{(}\sup_{s\in[0,t]}|\beta W_{s}|>x\Bigr{)}\leq 4\mathop{}\!\mathbb{P}(W_{t}>x/\beta)\leq 4\exp\Bigl{(}-\frac{x^{2}}{2\beta t}\Bigr{)}\leq\frac{4}{\mathrm{e}}\frac{2t\beta}{x^{2}}.$
A union bound then yields:
$\mathop{}\!\mathbb{P}\Bigl{(}\sup_{s\in[0,t]}|X_{s}|>x\Bigr{)}+t\,\overline{\pi}(x)$ $\leq\mathop{}\!\mathbb{P}\Bigl{(}\sup_{s\in[0,t]}|\beta W_{s}|>x\Bigr{)}+\mathop{}\!\mathbb{P}\Bigl{(}\sup_{s\in[0,t]}|Y_{s}|>x\Bigr{)}+t\,\overline{\pi}(x)$
$\leq\frac{8}{\mathrm{e}}t\beta x^{-2}+C\,t\,h(x)+t\,\overline{\pi}(x),$
which is bounded by some universal constant times $t$ times
$x^{-2}\int_{(0,x]}r^{2}\pi(\mathop{}\!\mathrm{d}r)+x^{-1}\int_{(x,\infty)}r\pi(\mathop{}\!\mathrm{d}r)+|d|x^{-1}+\beta x^{-2}.$
Let $\psi_{0}(\lambda)=\psi(\lambda)+d\lambda-\beta\lambda^{2}$ denote the Laplace exponent of the process with its drift and Gaussian component removed and let $\phi=\psi_{0}^{\prime}$ denote its derivative. Then $\phi$ is the Laplace exponent of a subordinator without drift and Lévy measure $\nu(\mathop{}\!\mathrm{d}r)=r\pi(\mathop{}\!\mathrm{d}r)$. In the last display, the sum of the first two terms, with the integrals, corresponds to $x^{-2}J_{\phi}(x)$ in the notation of [Duq12, Equation 47]. The next equation there, taken from [Ber96, Proposition III.1], shows that it lies between two universal constant times $x^{-1}\phi(x^{-1})=x^{-1}\psi_{0}^{\prime}(x^{-1})$, which itself, by [Duq12, Equation 50] lies between two universal constant times $\psi_{0}(x^{-1})=\psi(x^{-1})+dx^{-1}-\beta x^{-2}$. Therefore, our last display lies between two universal constant times $\psi(x^{-1})+(d+|d|)x^{-1}=\psi(x^{-1})+2d^{+}x^{-1}$. ∎
The exact result that we shall need in order to prove Proposition 5.1, based on this lemma, is the following. Recall that $\boldsymbol{\eta}$ is the exponent defined in (1.2). For any $\kappa,t>0$, we let $N_{\kappa}(t)=\#\{s\in[0,t]\colon\Delta X_{s}>\kappa^{-1}\}$ and $X^{\kappa}=(X^{\kappa}_{t})_{t\geq 0}$ the process obtained from $X$ by removing all these jumps larger than $\kappa^{-1}$.
Corollary 5.3.
Fix $N\geq 1$. There exists a constant $C_{N}$ such that for any $t>0$ and any $\delta>0$, we have
$\mathop{}\!\mathbb{P}\Bigl{(}\sup_{s\in[0,t]}|X^{\kappa}_{s}|>2N\kappa^{-1}\Bigr{)}+\mathop{}\!\mathbb{P}(N_{\kappa}(t)\geq N)\leq C_{N}t^{N}\kappa^{N(\boldsymbol{\eta}+\delta)},$
for all $\kappa>0$ large enough, depending in $\delta$ and $\psi$.
@@
Note that $X^{\kappa}$ is a Lévy process with the same drift and Gaussian component as $X$ and with Lévy measure $\pi^{\kappa}(\cdot)=\pi(\cdot\cap[0,\kappa^{-1}])$. Since its Laplace exponent is smaller than or equal to $\psi$, then we may apply Lemma 5.2 to $X^{\kappa}$ with $\psi$ in the upper bound. Because $X^{\kappa}$ makes only jumps smaller than or equal to $\kappa^{-1}$, at the first time it exceeds $x$ in absolute value, it cannot exceed $x+\kappa^{-1}$. With the help of the strong Markov property, we obtain by induction:
$\mathop{}\!\mathbb{P}\Bigl{(}\sup_{s\in[0,t]}|X^{\kappa}_{s}|>2N\kappa^{-1}\Bigr{)}\leq\bigl{(}C\,t(2d^{+}\kappa+\psi(\kappa))\bigr{)}^{N}.$
In addition, the random variable $N_{\kappa}(t)$ has the Poisson distribution with rate $t\,\overline{\pi}(\kappa^{-1})$. The Chernoff bound shows that the tail probability of a Poisson random variable $Y$ with rate $\lambda$ satisfies $\mathop{}\!\mathbb{P}(Y\geq k)\leq\operatorname{e}^{-\lambda}(\operatorname{e}\lambda/k)^{k}$ for every $k>\lambda$. Thus, for $c_{N}=(\operatorname{e}/N)^{N}$, we have
$\mathop{}\!\mathbb{P}(N_{\kappa}(t)\geq N)\leq c_{N}\bigl{(}t\,\overline{\pi}(\kappa^{-1})\bigr{)}^{N}\leq c_{N}\bigl{(}C\,t(2d^{+}\kappa+\psi(\kappa))\bigr{)}^{N}$
by Lemma 5.2. By the very definition of $\boldsymbol{\eta}$ we have $\psi(\kappa)=o(\kappa^{\boldsymbol{\eta}+\delta})$ as $\kappa\to\infty$ as well as $\kappa=o(\kappa^{\boldsymbol{\eta}+\delta})$ since $\boldsymbol{\eta}+\delta>1$, which concludes the proof. ∎
We may now prove our Hölder continuity result.
Proof of Proposition 5.1.
@@
For $1\leq n<n_{0}$, there are only finitely many quantities involved, so we infer that almost surely, there exists a now random $C_{a}\in(0,\infty)$ such that (5.1) holds for every pair $(s,t)$ of the form $(i2^{-n},(i+1)2^{-n})$ with $n\geq 1$ and $i\in\{0,\dots,2^{n}-1\}$. It then extends to all dyadic numbers by a routine argument, after multiplying $C_{a}$ by $2/(1-2^{-a})$, and then to the whole interval $[0,1]$ by density, using that $d_{\mathscr{L}(X^{\mathrm{ex}})}$ is continuous almost surely.
It remains to prove (5.2). Let us first replace $X^{\mathrm{ex}}$ by the unconditioned process $X$ and then argue by absolute continuity. According to (2.7), almost surely, for $s<t$, we have:
$d_{\mathscr{L}(X)}(s,t)\leq X_{s}+X_{t-}-2\inf_{r\in[s,t)}X_{r}\leq 4\sup_{r\in[s,t)}|X_{r}-X_{s}|.$
@@
Recall indeed that, informally the term in the middle is the length of the path if one goes from $s$ to $t$ in the looptree by always following a cycle on its right, instead of optimising between left and right. Note that the value of the jump of $X$ at time $t$, if any, does not affect the looptree distance, and not even the upper bound. This bound is too crude to be used directly because of the jumps of $X$, but we shall use it between consecutive large jumps since the possible jumps at time $s$ and $t$ do not appear in the upper bound.
@@
Precisely, fix $\delta>0$, let $\theta=(1+\delta)(\boldsymbol{\eta}+\delta)$, and let us prove (5.2) with $a=1/\theta$ by relying on Lemma 5.2 and Corollary 5.3 with $\kappa=2^{n/\theta}$. Indeed, fix $n\geq 1$ and $i\in\{0,\dots,2^{n}-1\}$. Let $t_{n,i}(1)<\dots<t_{n,i}(N_{n,i})$ denote the increasing enumeration of the set of times $\{t\in(i2^{-n},(i+1)2^{-n})\colon\Delta X_{t}>2^{-n/\theta}\}$. Set also $t_{n,i}(0)=i2^{-n}$ and $t_{n,i}(N_{n,i}+1)=(i+1)2^{-n}$. Then combining the last display with the triangle inequality, we have:
$d_{\mathscr{L}(X)}(i2^{-n},(i+1)2^{-n})\leq 4\sum_{j=0}^{N_{n,i}}\sup_{r\in[t_{n,i}(j),t_{n,i}(j+1))}|X_{r}-X_{t_{n,i}(j)}|.$ (5.3)
Note that conditionally on the $t_{n,i}(j)$’s, the paths between two consecutive such times behave like the process $X^{\kappa}$ with $\kappa=2^{n/\theta}$ from Corollary 5.3 whereas $N_{n,i}$ has the same law as $N_{2^{n/\theta}}(2^{-n})$. Some union bounds, invariance by time translation of the Lévy process, and the upper bound $t_{n,i}(j+1)-t_{n,i}(j)\leq 2^{-n}$, lead to the following bound for every $N\geq 1$:
$\mathop{}\!\mathbb{P}\Bigl{(}\sum_{j=0}^{N_{n,i}}\sup_{r\in[t_{n,i}(j),t_{n,i}(j+1)]}|X_{r}-X_{t_{n,i}(j)}|>2N^{2}2^{-n/\theta}\Bigr{)}$
$\leq\mathop{}\!\mathbb{P}(N_{n,i}\geq N)+\mathop{}\!\mathbb{P}\Bigl{(}N_{n,i}<N\enskip\text{and}\enskip\sup_{0\leq j<N}\sup_{r\in[t_{n,i}(j),t_{n,i}(j+1)]}|X_{r}-X_{t_{n,i}(j)}|>2N2^{-n/\theta}\Bigr{)}$
$\leq\mathop{}\!\mathbb{P}(N_{2^{n/\theta}}(2^{-n})\geq N)+N\mathop{}\!\mathbb{P}\Bigl{(}\sup_{r\in[0,2^{-n}]}|X^{2^{n/\theta}}_{r}|>2N2^{-n/\theta}\Bigr{)}.$
According to Corollary 5.3, for every $n$ large enough, the last line is bounded above by some constant that only depends on $N$ times
$(2^{-n}2^{n(\boldsymbol{\eta}+\delta)/\theta})^{N}=2^{-nN\delta/(1+\delta)}.$
Choose any $N>(1+\delta)/\delta$ and let $q=\delta N/(1+\delta)-1>0$. Then a last union bound implies that, for some constant $C_{N}$, for every $n$ large enough,
$\mathop{}\!\mathbb{P}\Bigl{(}\exists 0\leq i<2^{n}\colon\sum_{j=0}^{N_{n,i}}\sup_{r\in[t_{n,i}(j),t_{n,i}(j+1)]}|X_{r}-X_{t_{n,i}(j)}|>2N^{2}2^{-n/\theta}\Bigr{)}\leq C_{N}2^{-nq},$
The Borel–Cantelli lemma shows that almost surely these events eventually fail for $n$ large enough.
@@
For the conditional law, recall that the first half the bridge $(X^{\mathrm{br}}_{t};0\leq t\leq 1/2)$ is absolutely continuous with respect to the unconditioned process $(X_{t};0\leq t\leq 1/2)$. Hence almost surely, these events eventually fail for $n$ large enough in $X^{\mathrm{br}}$, when restricting to $i\in\{0,\dots,2^{n-1}\}$. In addition, the time- and space-reversal of the second half of the bridge has the same law as the first half, so this also holds for the second half and thus for the entire bridge. Finally we transfer it to the excursion by applying the Vervaat transform. Notice that the latter changes the starting point and thus shifts the intervals, but this only mildly affects the constant $2N^{2}$.
In each case, we obtain the desired Hölder property (5.2) thanks to the deterministic bound (5.3). ∎
@@
If one wanted to replace the exponent $1/\boldsymbol{\eta}$ by the smaller one $1/\boldsymbol{\gamma}$, then instead of the upper bound $\psi(x)\leq x^{\boldsymbol{\eta}+\delta}$ for every $x$ large enough, one would only have the existence of a sequence $x_{n}\to\infty$ along which $\psi(x_{n})\leq x_{n}^{\boldsymbol{\gamma}+\delta}$. Taking a subsequence if necessary, we may assume that $\sum_{n}x_{n}^{-1}<\infty$ and then the preceding argument shows that instead of (5.2) for any $a\in(0,1/\boldsymbol{\gamma})$, almost surely there exist $K_{a}>0$ and a sequence $\varepsilon_{n}\downarrow 0$ such that, for every $n$ large enough,
@@
We then turn to the map $M$ and the label process $Z^{\mathrm{ex}}$. Informally, we simply lose a factor $1/2$ due to the regularity of the Brownian labels. Again, we state the result under the excursion law but it holds similarly under the unconditioned or bridge law.